OpenStudy (anonymous):

find the zeros of each function by using a graph and a table? F(X)=x^2 +4x-5

OpenStudy (tkhunny):

You should start building your table. I would suggest looking somewhere between x = -5 and x = +2

OpenStudy (anonymous):

i dont even know what this means

OpenStudy (tkhunny):

You have no idea how to build a table of value?

OpenStudy (anonymous):

i dont know how to do this type of problemn and honestly no im illiterate to all math where do i get the points from

OpenStudy (tkhunny):

You are in the wrong class. You should go have a talk with your academic adviser and get into the right class. Given f(x)=x^2 +4x-5, We have f(0) = 0^2 + 4(0) -5 = 0 + 0 - 5 = -5 We have f(1) = 1^2 + 4(1) -5 = 1 + 4 - 5 = 0 <== Whoops! Accidentally found one. Keep building the table. Find the other one.

OpenStudy (anonymous):

Thank you!